### v2

# Temporal Feature Learning in Weakly Labelled Bioacoustic Cetacean Datasets via a Variational Autoencoder and Temporal Convolutional Network: An Interdisciplinary Approach

Abstract
Bioacoustics data from Passive acoustic monitoring (PAM) poses a unique set of challenges for classification, particularly the limited availability of complete and reliable labels in datasets due to annotation uncertainty, biological complexity due the heterogeneity in duration of cetacean vocalizations, and masking of target sounds due to environmental and anthropogenic noise. This means that data is often weakly labelled, with annotations indicating presence/absence of species over several minutes. In order to effectively capture the complex temporal patterns and key features of lengthy continuous audio segments, we propose an interdisciplinary framework comprising dataset standardisation, feature extraction via Variational Autoencoders (VAE) and classification via Temporal Convolutional Networks (TCN). This approach eliminates the necessity for manual threshold setting or time-consuming strong labelling. To demonstrate the effectiveness of our approach, we use sperm whale (Physeter macrocephalus) click trains in 4-minute recordings as a case study, from a dataset comprising diverse sources and deployment conditions to maximise generalisability. The value of feature extraction via the VAE is demonstrated by comparing classification performance against the traditional and explainable approach of expert handpicking of features. The TCN demonstrated robust classification capabilities achieving AUC scores exceeding 0.9.
I Introduction
Passive Acoustic Monitoring (PAM) is a method used to survey and monitor wildlife and environments using acoustic recorders. In the case of marine animals, PAM offers the advantage of enabling continuous data collection, even under adverse weather conditions or when visual surveys are unfeasible. As our capacity for collecting large quantities of acoustic data continues to grow, so does the demand for automated detection systems to effectively process and analyse it. While PAM has traditionally relied upon many different computational techniques for data analysis [32, 18, 11], Deep Learning (DL) methods have gained popularity for tasks that require detection of animal sounds from large acoustic datasets [44].
However, underwater PAM datasets pose a unique set of challenges when it comes to training DL methods, arising from the substantial cost associated with data collection and auditing, often aggravated by the scarcity of target events in the datasets. This results in considerably small annotated datasets with few instances of each signal of interest, both orders of magnitude below the dimensions of typical benchmark datasets [48, 53, 24]. Moreover, bioacoustics baseline datasets are often derived from limited surveys using single equipment types and narrow spatiotemporal ranges [8, 9, 10]. While these datasets are very valuable to establish a solid baseline for DL biacoustic studies, their reduced scope limits the generalisability of the models trained [36]. Finally, most underwater PAM data are annotated as presence/absence of vocalizations over periods of several minutes rather than individual calls, a practice known as weak labelling. Two common annotation approaches are auditing sampled intervals to mark species presence and marking acoustic event boundaries, which may include intermittent silence. These practices balance cost-effectiveness with field practicality, while also reducing variability between annotators [37]. However, weak labelling approaches significantly increase computational demands by generating data that is larger and more complex than that produced by strong labels. In turn, weak labelled datasets require larger and more complex models to process compared to finely annotated data, often rendering it practically infeasible.
One case for which weak labeling is especially fitting is sperm whale (Physeter macrocephalus) vocalizations. Sperm whales emit frequent and easily detectable clicks for the majority of their dive time. Maximum source levels of 240dB re 1 µPa have been recorded for these clicks [35], with click lengths of 1ms and average inter-click intervals of around 0.5 seconds between regular vocalisations [51]. The regularity of these clicks makes PAM a particularly suitable technique to study sperm whales, and because they are widely distributed, there are dozens of underwater datasets featuring their vocalizations. The acoustic and spectral properties of individual sperm whale clicks contain limited information to distinguish them from other transients, but their regularity in terms of amplitude and inter-click-intervals provides crucial contextual information. This makes weak labeling a practical choice for annotating sperm whale acoustic data, as it enables the annotator to leverage local information of individual clicks and broader temporal patterns into the decision. An effective detection system should therefore integrate both levels of spatiotemporal information to reliably identify click trains.
A suite of algorithms have been developed to detect sperm whale echolocation clicks, including approaches based on the time-frequency analysis of the clicks [33, 34], energy comparisons [22], or the use of the Teager-Kaiser Energy Operator [19]. In the field of DL, there has been work using convolutional neural networks to distinguish between spectrograms containing or not containing odontocete [30] or, more specifically, sperm whale [3] clicks. However, all such studies focus on detections of individual clicks, rather than incorporating multiple clicks into the decision process, limiting capture of the full vocalizations. Recent studies have attempted to address the task of detecting sperm whale click trains through heuristic methods, including the work of Macaulay [31] for detecting porpoise click trains, later applied to sperm whales in Webber et al. [49]. Another approach used deep learning models to classify long sequences (4 minutes) based on the sequence of click detections on shorter (5 second) windows [14]. While both approaches were successful in their respective tasks, the method proposed by Macaulay [31] requires extensive manual fine-tuning to suit specific datasets, and the DL approach of Garrobé Fonollosa et al. [14] is limited by the requirement for finely labeled data (in this case, annotations at a 5-second level), which is not always available.
This study proposes a new interdisciplinary methodology for detecting sperm whale click trains that addresses two of the most prevalent challenges of working with the complexities of bioacoustics data and training DL models for acoustic detection. Firstly, we tackle the issue of small dataset sizes and high variability in environmental and anthropological noise and acoustic data across different sources and geographical regions by curating a dataset from multiple sources collected under various deployment conditions. Secondly, we address the problem of weak labeling by proposing a workflow that leverages existing weakly annotated data to train robust DL classification models. This comprises feature extraction via an unsupervised variational autoencoder (VAE) [20] that captures complex and temporal patterns, both locally and long-term with its generative nature allowing it to learn robust features from noisy and complex time-series data. This avoids the need for setting manual thresholds or providing strong labels. The extracted features are then classified via a Temporal Convolutional Network (TCN), a DL architecture tailored for sequence processing [1], that have been successfully employed for audio data classification in recent years [1, 7, 54] to capture long-term dependencies and temporal context. To the best of our knowledge, this paper is the first to propose the VAE-TCN framework in the context of bioacoustic data detection and classification.
Traditional bioacoustic detection often relies on hand-made parameters such as peak frequencies, spectral shape, and inter-click intervals [46, 45, 33]. While recent methods using energy-temporal features [27] or acoustic indices [12] show promise, they often fail in new soundscapes [42]. Autoencoders [5] and VAEs have been shown to surpass handcrafted features for tasks such as species discrimination [16, 40] and call clustering [38], but direct comparisons for complex tasks remain lacking. We assessed the ability of VAEs to learn meaningful patterns and extract information from waveforms and spectrogram representations without prior knowledge of the target sounds. In order to evaluate the learning efficacy of VAEs, we compare them to traditional, expert-led methods used in earlier works [6] to select acoustic parameters, including root-mean-square calculations, energy level sums, and spectral properties.
The contributions of this study can be summarized as follows:
  1. 1.
    Curation of PAM recordings from various sources and deployment conditions into a comprehensive dataset that includes different types of background noise and other vocalising species to promote generalisation of the solution. Recordings from different studies were re-sampled and cleaned, with labels standardised and erroneous annotations removed. This dataset is weakly labelled with annotations for periods of multiple minutes, and containing well-balanced instances of the signal of interest (i.e., sperm whale vocalizations) and anthropogenic and environmental noise.
  2. 2.
    Feature extraction via VAEs to learn acoustic features from audio segments or click trains without requiring additional, time-consuming and thus costly, labelling.
  3. 3.
    Classification on 4-min long vocalization sequences, via TCNs trained on the above VAE-extracted features.
  4. 4.
    Evaluation of the efficacy of the extracted features in terms of classification accuracy against conventionally employed handcrafted acoustic features.
  5. 5.
    Evaluating the importance of annotation length on the different models’ performance to determine whether a greater emphasis on temporal context, or higher temporal resolution in extracted features, influences performance.
II Methods
Our proposed framework follows a systematic workflow comprising four stages, as shown in FIG.  1. Each of the four stages is described next.
Figure 1: Schematic diagram outlining the methodology employed to develop an acoustic classification framework from diverse datasets of weakly labelled data. The four main steps followed for this study are (1) standardising annotations and recordings from different sources and creating two datasets that were representative of the variabilities in anthropogenic and environmental noise, (2) VAE-based feature extraction, (3) detection and classification of sperm whale click trains based on TCN, and finally (4), evaluation of the value of feature extraction vs handpicked, expert-led features and annotation length of the temporal sequence.
II.1 PAM dataset curation
To enhance the robustness and generalisability of the solution, a comprehensive dataset was compiled from data collected by multiple studies utilising various sensors and recording protocols in different types of environment, covering different geographical and temporal scales (Table 1). By collating data from various sources, we aim to ensure robustness to the variabilities in the PAM datasets across various factors that would otherwise impact the classifier’s generalisation ability. These factors included:
  1. 1.
    Variations in the rhythmic patterns of sperm whale vocalisations, influenced by factors such as the number of sperm whales present and their behavioural state.
  2. 2.
    Diverse recording conditions and deployment specifics, such as the recording depth, proximity of the surface or seafloor, or frequency response of the recording systems.
  3. 3.
    The presence of anthropogenic and natural ambient sounds, as well as other vocalising species, and signal-to-noise ratio of the recordings.
  4. 4.
    Annotation variability from individual-level click labelling to 4-minute click train labelling to semi-automatic labelling.
Table 1: Geographical location, temporal scale, recording set-up and labelling process used in each of the six surveys that provided data to the present study.
Dataset name Location Time period Recording setup Labelling process
AS Atlantic, Southern, Arctic and Indian Oceans 2019 - 2021 Towed hydrophone Individual clicks labelled using PAMGuard automatic detector, with all detections reviewed by human analyst.
BAL_1 Balearic Sea 2015 - 2018 Ecological Acoustic Recorders (Oceanwide Science) sampling at 64 kHz for 4 min. every 30. 4-min recordings labelled as either containing sperm whale click trains, possible single vocalisations, or no sperm whale sounds.
BAL_2 2015 - 2017
BAL_3 2018 - 2019 SoundTraps (Ocean Instruments NZ) sampling at 96 kHz for 4 min. every 30.
CAL Southern California Blight 2009 - 2013 High-frequency acoustic packages [52] sampling at 100kHz - 160 kHz. Depth $\sim$ 65m-1300m Start and end time of odontocete encounters annotated.
CS Caribbean Sea 2019 - 2020 Custom built towed hydrophone sampling continuously at 96 kHz. Depth $\sim$ 10m. 4min every 30 audited in the field. Cetaceans and anthropogenic sounds annotated for the whole period.
ICE Iceland 2021 - 2022 Stereo deep water recorders (Loggerhead Instruments LS2X) at depth $\sim$ 2300m sampling at 96kHz for 5 minutes every 15. Files deemed positive by peak-to-peak SPL detector were manually audited and annotated. All clicks in each file were assigned the same label.
MED Mediterranean Sea 2004 - 2005 Towed array sampling at 48 kHz. Depth $\sim$ 10m. Individual clicks manually annotated using rainbowclick software.
The main dataset used in this study comprises data from six distinct research studies conducted across various global locations, with data collection efforts in Atlantic, Southern, Arctic and Indian oceans (AS dataset, Webber et al. [49]), the Balearic Sea (BAL datasets, Garrobé Fonollosa et al. [14]), the California Blight (CAL dataset, DCLDE [8]), the Caribbean Sea (CS dataset, Vachon et al. [47]), the Icelandic off-shore waters (ICE dataset, Wensveen et al. [50]) and the Mediterranean Sea (MED dataset, Lewis et al. [26]). The BAL dataset was divided into its 3 deployments (BAL_1, BAL_2, BAL_3) given that there was enough data available and previous studies [15] suggest high variability between deployments. The AS, CS and MED datasets were collected using towed hydrophones, while the BAL, CAL, CS and ICE data were collected using moored autonomous recorders. The detail and quality of the annotations also varied greatly from dataset to dataset, going from individual-level click labelling (MED dataset), to labelling at a 4-minute level (BAL and CS datasets), or semi-automatic labelling using automated acoustic algorithms to do an initial processing of the data (AS and ICE datasets).
Given the diverse recording conditions and labelling formats across data from different deployments, a necessary first step was standardising both recordings and labels. A uniform labelling format was established, which entailed dividing recordings into 4-minute segments that were then labelled either as positive (containing sperm whale clicks) or negative (not containing sperm whale clicks), derived from the original annotations. The 4-minute length was chosen due to it being the finest common resolution available across all sources. Additional anthropological and environmental information, such as the presence of other sound sources like ships and other vocalising animals, was retained when available. These annotations were solely kept as references to ensure a diverse range of anthropological and environmental noise sources in the final dataset but were not used as labels for training.
Imprecisely labelled files (those labelled as containing click trains from unknown animals or unknown odontocetes) were excluded from the training set, as were files labelled as containing single sperm whale clicks but not click trains, due to the variability in this category between expert analysts [14]. The final dataset comprised all available positive files along with a randomly chosen set of negative files, including files labelled as containing other sources of impulsive noise and those without any marked sources, ensuring a 50/50 split between positive and negative files.
In addition to the above 4-minute segment dataset, a dataset comprising 30-second audio files was constructed that served as the basis for examining the trade-off between increased temporal context for decision-making and shorter audio segments allowing for higher resolution features to be generated and extracted. This dataset was comprised of data from the AS and MED datasets, the two datasets where annotations were available at a single click level (Table 1). The same pre-processing steps as in the 4-minute dataset were applied to the 30-second segments.
II.2 Pre-processing
After selecting the 4-minute and 30-second segments and standardising the labels to either positive or negative for the whole segment, a pre-processing stage was implemented to standardise the recordings across diverse deployments. In cases of recordings featuring multiple channels, only the first channel was retained. Furthermore, any DC offset present in the signal was removed. Subsequently, all files were decimated to a uniform sampling rate of 48 kHz. To eliminate low-frequency noise, the signal was filtered via a 1-20kHz bandpass 6-pole Butterworth filter, leaving a frequency range that captures the majority of the sperm whale click energy to analyse [17].
An impulsive noise detector algorithm was used to identify and select up to 10 transient sounds with the highest amplitude within each 4 minute segment, regardless of what the recording had been labeled as. Transient sounds were detected using a single pole filter with a 6dB threshold (see Alg. 1 shown in FIG.  2). Subsequently, waveform segments of varying lengths were extracted from these transients to train different feature engineering methods, ensuring the random placement of amplitude peaks within each segment. Additionally,5 randomly sampled segments per file where no transients was detected were included in order to incorporate examples of background noise into the optimisation of feature engineering methods. The proposed approach acknowledges that, while impulsive noises detected in a 4-min sequence might not necessarily share the label of the parent audio file, a considerable proportion are expected to do so, such that this procedure would produce a dataset representative of the different sound sources present in the data. In summary, for each 4-min labelled segment, we extract up to 10 top high amplitude sequences plus 5 randomly selected quiet periods to train the feature extractors.
Figure 2: Impulsive noise detector algorithm
For audio classification at the file level, recordings were divided into shorter, non-overlapping windows of two sizes: 512 and 2048 samples (0.01s and 0.04s, respectively). The two window lengths were used to investigate the trade-off between temporal resolution and sequence length. This process generated 22,500 and 5,625 windows, respectively, for 4-minute files, and 2,812 and 703 windows for 30-second files (Table 7). The feature vectors detailed in the previous section were concatenated, transforming each audio file into a sequence of size ($m$, $n$), where $m$ represents the number of extracted features and $n$ denotes the number of non-overlapping windows in the file. Each of the $m$ features, both from handcrafted features and VAE embeddings, was normalised using median and standard deviations of the feature on the values obtained on the short audio segment dataset.
II.3 Feature engineering
Features were generated two-fold: (i) expert-based or handcrafted feature selection and (ii) feature extraction via variational autoencoders (VAEs). Whilst the former in more transparent and rooted in marine science [6], it may not necessarily be optimised for the classifier. Feature extraction is not explainable to a human but tends to learn compressed, probabilistic latent representations of the PAM segments, capturing key temporal dynamics and structure in the click trains.
II.3.1 Handcrafted feature generation
Several acoustic features were selected from the non-overlapping short audio segments generated as described in Subsection II.2, namely the root-mean-square of the time-series for a specific frequency band $f_{1},f_{2}]$ ($RMS_{f_{1},f_{2}}$), peak and mean frequencies, spectral width, and energy sum over three frequency bands (1kHz - 4kHz, 4kHz-8kHz, 8kHz-16kHz) (see Figure 3 for an example and Table 2 for formal definition of each handcrafted feature). The choice of these parameters was motivated by earlier works on the detection-classification of odontocete vocalisations [6], and the frequency bands were chosen to correspond to the distinct bandwidths in the sperm whale click spectral profile [17].
Figure 3: Parametrisation of a sperm whale click waveform and spectrum using the handpicked features described in Table 2.
Table 2: Handcrafted features
Feature Formula Parameters
Root mean square $RMS_{f_{1},f_{2}}=\sqrt{\frac{1}{n}\displaystyle\sum_{i=1}^{n}x\left[f_{1},f_{2}\right]_{i}^{2}}$ $n$: number of samples in window.
$x\left[f_{1},f_{2}\right]$: waveform filtered between the frequencies of $f_{1}$ and $f_{2}$.
Peak frequency $pkf=argmax\left(s(f)\right)$ $f$: vector of frequencies returned by fft
$s$: squared smoothed spectral vector (from fft)
Mean frequency $\overline{f}=\frac{s\cdot f}{\sum s}$ $f$: vector of frequencies returned by fft. Frequency range 1kHz-20kHz.
$s$: squared smoothed spectral vector (from fft)
Energy sum $E_{f_{1},f_{2}}=\displaystyle\sum_{f\in[f_{1},f_{2})}s[f]$ $s[f]$: squared smoothed spectral vector at frequency $f$ (from fft)
Spectral width $\omega=f_{M}-f_{m}$ $f_{M}$ and $f_{m}$: maximum and minimum frequencies with an amplitude above the peak amplitude minus 8dB and with no more than 3 frequency values in a row below that threshold in between the peak and those frequencies.
The extracted handpicked features were combined into different sets of parameters (Table 3). These combinations include: RMS only vectors over 1, 3, and 5 frequency bands, which were chosen for their ability to capture amplitude peaks across different frequency ranges; a set including 1 RMS band, the spectral features and the energy values, selected to incorporate amplitude, energetic and spectral information; and a subset of the previous one including only one RMS value along with peak frequency, mean frequency, and spectral width, which we chose to test whether dividing the signal into energy bandwidths was being used or simply providing redundant information.
Table 3: Combinations of handpicked acoustic features used in the study
Type Number of parameters Parameters
RMS only 1 $RMS_{1-20kHz}$
3 $RMS_{1-6kHz}$, $RMS_{6-12kHz}$, $RMS_{12-20kHz}$
5 $RMS_{1-2kHz}$, $RMS_{2-4kHz}$, $RMS_{4-8kHz}$, $RMS_{8-16kHz}$, $RMS_{16-20kHz}$,
Spectral features 7 $RMS_{1-20kHz}$, pkfpkf, $\overline{f}$, $\omega$, $E_{1-4kHz}$, $E_{4-8kHz}$, $E_{8-16kHz}$
II.3.2 Feature Extraction with Variational AutoEncoders (VAE)
VAEs were trained on three different forms of input data: waveforms, spectral profiles, and spectrogram representations. To generate the spectrograms and spectral profiles we used a 512-point Hann window fft with a 50% overlap. Spectral profiles were computed on waveforms of length 512 samples, and frequencies below 1kHz or above 20kHz were discarded, leading to a final sequence of length 200, which was then smoothed using an 8-point moving average. Additionally, longer spectrograms were computed for waveform sections of length 32,768 samples (0.68 seconds), so as the resulting spectrogram to be 128 FTTs in length giving an image of width 128 pixels. Again, frequencies below 1kHz and above 20kHz were discarded to center sperm whale clicks in the spectrogram [17], and the final representation was resized to be of size $\times128$ pixels.
A 2D-VAE that utilizes a ResNet-18 architecture [43], a network that has consistently exhibited good performance in bioacoustics call detection literature [2, 21, 28, 41] was tuned for extracting features from spectograms. On the other hand, waveforms and smoothed spectral profiles both were autoencoded using a 1D-VAE [29]. Each model was trained for different latent sizes: 24, 32, 48 and 64. The 2-D VAE was implemented for two additional latent sizes, 96 and 128. Since VAEs return the parameters of a normal distribution in the latent space, for each input data point encoded into a space of dimension $n$, a vector of size 2nn is outputted where the first half of the vector corresponds to the means of the normal distribution and the second half to the variances. This strategy of mapping data to normal distribution parameters facilitates interpolation between data points, enhancing the robustness of embeddings against various forms of noise and domain shifts [39].
II.4 TCN classification
Figure 4: Schematic of the TCN architecture (left), a temporal block in the TCN (middle) and a dilated casual convolutional layer with a kernel size (k) of 3 and a dilation (d) of 2). The deep stack of dilated convolutions allows the TCN to capture long-range temporal patterns, making it a suitable architecture to detect the rythmic vocalisation patterns of sperm whales.
A TCN was used for classifying full recordings from the sequence of extracted features based on the presence or absence of sperm whale vocalisations. The architecture shown in Figure 4 was adopted due to its proven ability to capture long-term dependencies and timing patterns in acoustic signals [25], which would be key for detecting groups of sperm whales of varying sizes and distinguishing them from other sources of impulsive noise. The TCN was of length $n$ and had $m$ channels.
All TCNs were trained in a supervised manner using the Adam optimizer and negative log likelihood loss function. Hyperparameters, including the number of hidden layers, kernel size, dropout rate, learning rate, and number of levels, were optimized using a grid search strategy (Table 6). The best-performing combination consisted of a batch size of 8, 8 hidden layers of size 25, a kernel size of 20, a dropout rate of 0.4, and a learning rate of 0.001. The final architectures for both the 4-minute and 30-second segment models are summarized in Table 7.
III Results
The performances metrics of the different trained TCNs were compared on an unseen subset of the data reserved for testing. A random +:15 split was used to partition the data into training and testing sets, and that split was consistent across models trained on the same length of recordings. Performance evaluation was conducted based on recall ($TP/P$) and false-positive-rate (FPRFPR; $FP/N$), where $P$ and NN represent the total number of positive and negative files (as annotated), and TPTP and FNFN denote correct and incorrect predictions of the model from the files labelled as containing sperm whales by humans. The best model is one that minimizes FPRFPR while simultaneously maximizing recall.
III.1 Data segments for training
The final 4-minute training set contained 7,668 audio files (Table 4) coming from 8 different deployments, with a 50/50 split between files that were deemed as containing and not containing sperm whale clicks. This dataset included data from different sperm whale populations spanning 20 years, 8 different geographical locations, 7 different recording devices and over 8 other vocalising species that produce echolocation clicks, as well as other sources of impulsive noise. The 30-second dataset contained a total of 9,912 files from 2 different sources, AS and MED (Table 4).
Table 4: The number of recordings in the merged dataset labelled as containing only sperm whales (SW), sperm whales and other sources of impulsive noise, other sources of impulsive noise only, and recordings audited but nothing noted down in the final training set separated by source. For the Balearic data recordings were only marked for the presence of sperm whale vocalisations, so no information on other sound sources was available. The numbers outside of parenthesis refer to the 4-minute dataset, and parenthesis to the 30-second dataset.
Dataset SW only SW + Other Other only Nothing annotated Total
AS 149 (1900) 8 (26) 123 (1055) 234 (871) 514 (3852)
BAL_1 93 - - 93 186
BAL_2 143 - - 142 285
BAL_3 273 - - 273 546
CAL 173 1 114 60 348
CS 773 560 720 607 2,660
ICE 365 558 379 544 1,846
MED 626 (3,003) 16 (31) 47 (230) 594 (2,796) 1,283 (6,060)
Total 2,595 (4903) 1,143 (57) 1,383 (1285) 2,547 (3667) 7,668 (9912)
A total of 101,210 short segments were identified to train the feature engineering methods to train the feature extraction methods. 71,893 of these were sampled using an impulsive noise detector. Out of these, 37,638 came from recordings labelled as containing sperm whale clicks and 34,206 came from recordings labelled as not containing sperm whale clicks. The remaining 29,317 short audio segments were randomly sampled from the long audio segments to incorporate a representation of background noise. It is important to note that these latter segments were sampled independently from the impulsive noise detections.
III.2 Feature extraction methods and training of 4-min audio file classifier
A total of 24 TCN models with optimised parameters and structure (Table 5) were trained on the collated 4-minute audio dataset for different forms of feature extraction (Figure 5). The best performance both in terms of high recall and low FPR is obtained using a 2D-VAE on spectrogram representations, followed by multi-channel RMS measures and a 1D-VAE on spectral profile. The performance of TCNs applied to encoded representations of waveforms yields the worst results every time, suggesting that the data inputted into the VAE was too noisy and variable to extract the meaningful information needed for this task.
Figure 5: Recall - false positive rate (FPR) relation on the validation set of 4-minute dataset for TCNs trained on parameters extracted using VAEs and traditional methods for acoustic feature extraction. Number in parenthesis refers to the size of the window over which the feature extraction was performed. Size of the extracted parameters for each of the non-overlapping windows is displayed in text next to each point. The most efficient detectors will be in the lower right quadrant, which in this case are the ones that work on sequences of VAE embeddings of spectrograms.
Figure 6: Box plot displaying recall (orange) and FPR (grey) scores on validation data divided by deployment for the best performing model for 3 different feature extraction types. Dots displaying data for the 8 deployments (Table1) are displayed next to each box plot. The best models will maximize recall while minimizing FPR scores, and thus will have the grey and orange boxes clearly separated. The figure shows that the detector that manages to perform better across data from different deployments is the one that works on sequences of embeddings of 2D spectrograms.
When looking at performance variability across sources, we see that the TCN trained on 2D-VAE embeddings from spectrogram data also sees the highest lowest Recall and the second lowest highest FPR (Figure 6). This suggests that this feature extraction method not only performs better overall but also shows a more stable performance across data from different sources.
III.3 The effect of annotation length
On 30-second recordings, the results demonstrate a notable improvement in both higher recalls and lower false positive rates for TCNs with different input features when compared to the features from the 4-minute recordings (Figure 7, Figure 8). The only two exceptions to this trend are TCNs trained on VAE embeddings from waveforms, which exhibit no improvement and are consistently worse-performing method, and TCNs trained on VAE embeddings from spectrograms, which only show a marginal performance improvement. Notably, for 30-second recording classification, TCNs trained on VAE embeddings for spectral profile show the most improvement and achieve the best overall performance across all tested methods for feature extraction.
Figure 7: Recall - false positive rate (FPR) relation on the validation set of 30-second files for TCNs trained on parameters extracted using VAEs and traditional methods for acoustic feature extraction. Number in parenthesis refers to the size of the window over which the feature extraction was performed. Size of the extracted parameters for each of the non-overlapping windows is displayed in text next to each point. The most efficient detectors will be in the lower right quadrant, which in this case are the ones that work on sequences of RMS values over 5 frequency bands and the ones that work on sequences of VAE embeddings of spectral profiles.
IV Discussion
In this study, we developed a novel interdisciplinary workflow designed to effectively utilise weakly labelled acoustic datasets, a prevalent challenge in bioacoustics research. Our proposed approach consisted of four stages: (i) robust dataset curation from multiple geographical recordings to include various sources of environmental and anthropological noise, (ii) feature generation via expert-based feature selection and unsupervised feature extraction via VAEs, (iii) classification at the recording level using TCNs, which incorporated temporal context into the decision making. This approach diverges from conventional segmentation strategies, which risk overlooking crucial contextual cues inherent in the longer intervals at which weak labels are set. We found that with the proposed approach TCNs were able to classify recordings of 4-minutes with a performance comparable to those carried out by human analysts [14].
We explored the effectiveness of both handpicked acoustic indices and VAEs for unsupervised feature extraction. Results showed that both handpicked features and VAE embeddings were able to provide enough information for a TCN to obtain recall rates of over 0.83 while maintaining a false detection rate below 0.13. TCNs trained on VAE embeddings of spectrogram representations performed marginally better on the task of classifying 4-minute recordings. Our results also showed that handpicked acoustic features were more prone to overfitting to some datasets and their recall rates varied significantly by deployment, while VAE embeddings achieved a more consistent performance across data sources. This finding is particularly significant to the bioacoustics community as it highlights the capacity of VAEs to offer a consistent and reliable feature representation, regardless of variations in data sources and deployment conditions.
Our approach differs from previous sperm whale detection approaches through three key innovations. First, we employ TCNs to model long-range dependencies in acoustic sequences, enabling detection decisions at biologically-relevant timescales for this species (e.g., minute-level click trains) rather than isolated click-level analysis typical of traditional energy detectors [19, 34, 3]. Second, we achieve robust performance over these long timescales without the need for manually set thresholds [31, 49] or finely labeled data [14]. Third, we integrate VAEs as feature extractors, simultaneously improving computational efficiency in the classification step while preserving enough information to distinguish sperm whale vocalizations from ambient noise. This combination of large-scale temporal modeling, reduced annotation dependence, and efficient data compression addresses fundamental limitations in current passive acoustic monitoring pipelines.
To our knowledge, this study also presents the largest and most varied sperm whale acoustic dataset curated to date, combining recordings from six different studies spanning decades, diverse geographic regions, and multiple deployment conditions. This variability, which includes differences in background noise, vocalising species, and recording set ups, is critical to training robust, generalizable models. This dataset provides a reference point for future research on weakly labeled bioacoustic detection, addressing a key need for open, high-quality bioacoustics baseline datasets [13].
Differences in performance between data sources could stem from variations in noise levels and recording conditions [4, 36], but annotation protocols and survey designs may also play a significant role. For example, aurally auditing four minutes of single-channel data may lead to missed detections of fainter, less regular sperm whale clicks, whereas manually reviewing each click in isolation allows for more thorough annotation. Moreover, the purpose of the study that generated a particular dataset also influences the detectability of the animals in the acoustic recordings. For example, the CS dataset was collected during a dedicated sperm whale behavioral study, where a vessel actively tracked whales to study social interactions. Consequently, the recordings predominantly contain high-amplitude, well-defined clicks from nearby individuals, introducing a proximity bias in the vocalization examples. In contrast, datasets from static hydrophones or line surveys are more likely to include faint, distant clicks, which poses challenges for both human annotators and automated detectors. When comparing detector performance across datasets, reported metrics reflect both the algorithm’s efficacy and dataset-specific annotation biases. All results must therefore be interpreted within their recording and annotation context.
The present study also revealed that classifiers trained and tested on shorter duration recordings (30 seconds instead of 4 minutes) consistently exhibited superior performance. Although this outcome was anticipated, it should be noted that the shorter dataset comprised recordings from only two sources, potentially contributing to lower data variation. However, what is worth noting is that in this scenario TCNs trained on embeddings of spectral profiles outperformed those trained on the embeddings of 2D spectrograms, which were the best performing method for the 4-minute audio files. This suggests that the length of the final vector of acoustic features became a limiting factor in the classifier’s performance for 4-minute recordings. Consequently, methods that generate shorter sequences yielded better performances, even at the expense of time resolution. However, for classification at shorter time scales, these features extracted at a higher resolution significantly benefit the TCN step in the classification task.
V Conclusion
Our study suggests a way forward in leveraging numerous existing annotated bioacoustics datasets to train automatic classification models, effectively overcoming previous limitations associated with weak labels. Notably, our ability to train the model on a scale as large as 4 minutes means that marine scientists can readily apply this approach to nearly all available annotated data without additional auditing, enabling for the processing of vast amounts of data efficiently. This capability opens doors for researchers to analyze higher volumes of marine acoustic data. Beyond marine science, our proposed approach has potential applicability in other acoustic fields, including soundscape analysis, environmental monitoring, or healthcare. Future research should focus on transferability of the best-performing models on distinct unseen new sources. Additionally, research will be carried out on more fine-tuned hybrid human-AI feature selection [23] for improved feature learning, robustness and transparency.
Acknowledgements.
This work was funded by the Scottish Universities Partnership for Environmental Research (SUPER) Doctoral Training Partnership (DTP), between the Universities of St Andrews and Strathclyde. We would like to express our gratitude to Dr Kirsten Young, Thomas Webber, Asociación TURSIOPS, Dr Felicia Vachon, Dr Hal Whitehead, Dr Paul J. Wensveen, Michelle Dutro, Caroline Haas, the Marine Conservation Research/International Fund for Animal Welfare, and Greenpeace International for supplying the essential data for this research. Collection of the Iceland data set was made possible by the Rannís Icelandic Research Fund (number 207081).
AUTHOR DECLARATIONS
Conflicts of Interest
The authors have no conflicts to disclose.
Data availability
The acoustic datasets are available on request from the corresponding author, LFG. The data are not publicly available due to size restrictions.
Source code and trained models are openly available at github.com/laiagf/WeakDetector.
Supplementary materials
Table 5: Layers and parameters of the TCN used for recording classification, where NN is the number of non-overlapping windows in the file and $m$ the number of extracted features.
Layer Output shape Param #
Temporal Block 1 (k = 20, d = 1, dropout = 0.4) [-1, 25, N] m*500 + 25
12,525
25*(m+1)
Temporal Block 2 (k = 20, d = 2, dropout = 0.4) [-1, 25, N] 12,525
12,525
Temporal Block 3 (k = 20, d = 4, dropout = 0.4) [-1, 25, N] 12,525
12,525
Temporal Block 4 (k = 20, d = 8, dropout = 0.4) [-1, 25, N] 12,525
12,525
Temporal Block 5 (k = 20, d = 16, dropout = 0.4) [-1, 25, N] 12,525
12,525
Temporal Block 6 (k = 20, d = 32, dropout = 0.4) [-1, 25, N] 12,525
12,525
Temporal Block 7 (k = 20, d = 64, dropout = 0.4) [-1, 25, N] 12,525
12,525
Temporal Block 8 (k = 20, d = 128, dropout = 0.4) [-1, 25, N] 12,525
12,525
Fully Connected Layer [-1, 2] 52
Parameter Values
Batch size 4, 8, 10 12, 16
Size of hidden layers 8, 12, 15, 20, 25
Number of hidden layers 4, 8, 12, 16, 20
Kernel size 10, 15, 20, 25
Dropout 0.1, 0.2, 0.3, 0.4
Learning rate 0.1, 0.01, 0.001, 0.0001
Table 6: Values tested for the TCN hyperparameter optimization. Chosen values are in bold.
Table 7: Summary of input data and sizes for all TCN models trained and tested.
Feature extraction method N. channels ($m$) Feature extraction window size 4-min TCN length ($n$) 30-sec TCN length ($n$)
RMS 1, 3, 5 512 samples 22500 2812
2048 samples 5625 703
Spectral parameters 7 512 samples 22500 2812
2048 samples 5625 703
1D-VAE on waveform 8, 16, 24, 32 512 samples 22500 2812
2048 samples 5625 703
1D-VAE on spectral profile 8, 16, 24, 32 512 samples 22500 2812
2D-VAE on spectrogram 8, 16, 24, 32, 64, 128 32768 samples 352 44
Figure 8: Recall - false positive rate (FPR) relation on the valiadion set of TCNs trained on parameters extracted using VAEs and traditional methods for 30-second (orange) and 4-minute (grey) audio clips. Best performing methods will be in the bottom right of each plot. Notably, TCNs exhibit superior performance in classifying shorter clips across all feature extraction methods. The only exception to that are TCNs trained on VAE embedding representations of spectrograms, which demonstrate consistent performance for both audio lengths.